# Revisiting Randomization in Greedy Model Search

### v1

@@ A.3 Proof for Error Improvement
$\alpha_{p}^{\star}=\log\Big{(}\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\cdot\frac{A}{B}\Big{)},$ (63)
@@
$\alpha_{p}^{\star}=\log\Big{(}\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\cdot\frac{A}{B}\Big{)}\geq\log\Big{(}\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\Big{)}=\alpha_{0}^{\star}.$ (64)
Thus, substituting $\alpha_{p}^{\star}$ and $\alpha_{0}^{\star}$ into (62) and taking $\epsilon=0.05$ gives us the lower bound on expected training error gap (41) as
@@
$\mathbb{E}\Big{[}\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\min_{m}\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}\Big{]}$ $\geq\frac{1}{4}\frac{\big{(}\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big{)}^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}.\qed$
A.3.1 Proofs for Degrees of Freedom Results
@@
Note that for $k=1,2,\ldots,p$, we have
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}\hat{\beta}_{j,p}x_{j,p},\quad\hat{f}_{k,p}=\sum_{j=1}^{k}\hat{\beta}_{j,p}x_{j,p},$ (65)
which implies
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}(\hat{f}_{j,p}-\hat{f}_{j-1,p}).$
Since degrees of freedom is a linear operator over estimators, we have
$\operatorname{df}(\hat{f}_{k,m,p})=\sum_{j=1}^{p}w_{j}^{k,m,p}\Delta(j),$ (66)
where $\Delta(j)\coloneqq\operatorname{df}(\hat{f}_{j,p})-\operatorname{df}(\hat{f}_{j-1,p})$, for $j=1,2,\ldots,p$, forms a non-increasing sequence by the concavity assumption. We claim that whenever $m^{\prime}>m$, the sequence $\big{\{}w_{j}^{k,m^{\prime},p}\big{\}}_{j=1}^{p}$ majorizes the sequence $\big{\{}w_{j}^{k,m,p}\big{\}}_{j=1}^{p}$. This means that for each $l\geq 1$,
$\sum_{j=1}^{l}w_{j}^{k,m^{\prime},p}\geq\sum_{j=1}^{l}w_{j}^{k,m,p},$ (67)
@@
Assuming the majorization in (67) holds, the theorem follows directly from Theorem A.3 in [38] by using the facts that both $w_{j}^{k,m,p}$ and $\Delta(j)$ are non-increasing in $j$.
We return to proving the claim. Let $S_{k,m,p}$ denote the indices of the features selected by one inner loop of RGS with parameters $k$, $m$, and $p$. We may write
$\begin{split}\sum_{j=1}^{l}w_{j}^{k,m,p}&=\sum_{j=1}^{l}\mathbb{P}(j\in S_{k,m,p})\\ &=\mathbb{E}\left[\sum_{j=1}^{l}\mathbb{1}(j\in S_{k,m,p})\right]\\ &=\mathbb{E}\left[Z_{k,m,p}(l)\right],\end{split}$ (68)
where $Z_{k,m,p}(l)\coloneqq|S_{k,m,p}\cap[l]|$. Evidently, (67) will be established if we can produce a coupling between $S_{k,m,p}$ and $S_{k,m^{\prime},p}$ such that $Z_{k,m^{\prime}p}(l)\geq Z_{k,m,p}(l)$ for $l=1,2,\ldots,p$.
We construct such a coupling by running two instances of RGS concurrently, one with parameter choice $m$ and the other with parameter choice $m^{\prime}$. At iteration $k+1$ of the algorithm, we first draw a subset $B_{m}$ of size $m$ from $[p-k]$ uniformly at random. We then draw a further subset of $m^{\prime}-m$ elements from $[p-k]\backslash B_{m}$ uniformly at random and take their union with $B_{m}$ to form $B_{m^{\prime}}$. We then construct the candidate subset $\mathcal{V}_{k+1,m,p}$ by ordering $[p]\backslash S_{k,m,p}$ from smallest to largest and picking the elements at positions in $B_{m}$. To construct the candidate subset $\mathcal{V}_{k+1,m^{\prime},p}$, we do the same, but with $[p]\backslash S_{k,m^{\prime},p}$ and $B_{m^{\prime}}$ instead of $[p]\backslash S_{k,m,p}$ and $B_{m}$ respectively.
We now show the desired property inductively in $k$, with the base case $k=0$ trivial. Assume the inductive hypothesis. Let $b\coloneqq\min B_{m}$. Then one can show that the index selected at iteration $k+1$ (with parameter $m$) is equal to $l_{*}\coloneqq\min\left\{l\colon b\leq l-Z_{k,m,p}(l)\right\}$. We therefore have
$Z_{k+1,m,p}(l)=\begin{cases}Z_{k,m,p}(l)&\text{if}~{}l<l_{*}\\ Z_{k,m,p}(l)+1&\text{if}~{}l\geq l_{*}.\end{cases}$ (69)
Similarly, denoting $b^{\prime}\coloneqq\min B_{m^{\prime}}$ and letting $l_{*}^{\prime}\coloneqq\min\left\{l\colon b^{\prime}\leq l-Z_{k,m^{\prime},p}(l)\right\}$, we get
$Z_{k+1,m^{\prime},p}(l)=\begin{cases}Z_{k,m^{\prime},p}(l)&\text{if}~{}l<l_{*}^{\prime}\\ Z_{k,m^{\prime},p}(l)+1&\text{if}~{}l\geq l_{*}^{\prime}.\end{cases}$ (70)
To compare $Z_{k+1,m^{\prime},p}$ and $Z_{k,m,p}$, we further set $\tilde{l}\coloneqq\min\left\{l\colon b\leq l-Z_{k,m^{\prime},p}(l)\right\}$ and construct the function $\tilde{Z}$ via
$\tilde{Z}(l)=\begin{cases}Z_{k,m^{\prime},p}(l)&\text{if}~{}l<\tilde{l}\\ Z_{k,m^{\prime},p}(l)+1&\text{if}~{}l\geq\tilde{l}.\end{cases}$ (71)
@@
$\tilde{Z}(l)-Z_{k+1,m,p}(l)=\begin{cases}Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)&\text{if}~{}l<l_{*}\\ Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)-1&\text{if}~{}l_{*}\leq l<\tilde{l}\\ Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)&\text{if}~{}l\geq\tilde{l}.\end{cases}$ (72)
@@
$Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)=(Z_{k,m^{\prime},p}(l)-l)-(Z_{k,m,p}(l)-l).$ (73)
Because these functions take integral values, the only way $Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)-1$ can be negative is if the right hand side of (73) is equal to 0. However, by the way in which $l_{*}$ is defined, we would have
$l-Z_{k,m^{\prime},p}(l)=l-Z_{k,m,p}(l)\geq b,$ (74)
@@
Utilizing the majorization in (67), we now turn to prove Theorem 4.3.
Proof of Theorem 4.3.
According to the decomposition (6), it suffices to prove
$\sum_{j=1}^{p}\hat{\beta}^{2}_{j,p}\big{(}w_{j}^{k,m,p}\big{)}^{2}\leq\sum_{j=1}^{p}\hat{\beta}^{2}_{j,p}\big{(}w_{j}^{k,m^{\prime},p}\big{)}^{2}.$ (75)
Applying Proposition A.3.a from [38], where in their notation we take $p_{j}=\hat{\beta}^{2}_{j,p}$ and $g(x)=x^{2}$, noting that $g$ is an increasing convex function on the positive real numbers, it follows that (75) holds if for any $l\geq 1$,
$\sum_{j=1}^{l}\hat{\beta}^{2}_{j,p}w_{j}^{k,m,p}\leq\sum_{j=1}^{l}\hat{\beta}^{2}_{j,p}w_{j}^{k,m^{\prime},p}.$ (76)
Applying Theorem A.3 from [38] and using the fact that $\hat{\beta}^{2}_{j,p}$ is non-increasing in $j$, we conclude that (76) holds for any $l\geq 1$. ∎
@@
$\hat{\beta}_{i,p}-\hat{\beta}_{j,p}=(\beta_{i,p}-\beta_{j,p})+\langle\mathbf{x}_{\cdot i,p}-\mathbf{x}_{\cdot j,p},\boldsymbol{\varepsilon}\rangle.$ (77)
@@
$\mathbb{P}\left(\hat{\beta}_{i}-\hat{\beta}_{j}<0\right)\leq(4\pi\sigma^{2}/n)^{-1/2}\exp(-n/4\sigma^{2}).$ (78)
@@
$\mathbb{P}\left(\hat{\beta}_{i}+\hat{\beta}_{j}<0\right)\leq(4\pi\sigma^{2}/n)^{-1/2}\exp(-n/4\sigma^{2}).$ (79)
@@
$\begin{split}\mathbb{P}\left(|\hat{\beta}_{j,p}|>|\hat{\beta}_{i,p}|\right)&\leq\mathbb{P}\left(|\hat{\beta}_{j,p}|>\hat{\beta}_{i,p}\right)\\ &\leq\mathbb{P}\left(\hat{\beta}_{j,p}>\hat{\beta}_{i,p}\right)+\mathbb{P}\left(-\hat{\beta}_{j,p}<\hat{\beta}_{i,p}\right)\\ &\leq(\pi\sigma^{2}/n)^{-1/2}\exp(-n/4\sigma^{2}).\end{split}$ (80)
Let $\mathcal{A}$ be the event on which $\operatorname{rank}(|\hat{\beta}_{j,p}|\leq k$, for $j=1,\ldots,k$. We have
$\begin{split}\mathbb{P}\left(\mathcal{A}\right)&=\mathbb{P}\left(|\hat{\beta}_{i,p}|>|\hat{\beta}_{j,p}|,1\leq i\leq k<j\leq p\right)\\ &\geq 1-k(p-k)(\pi\sigma^{2}/n)^{-1/2}\exp(-n/4\sigma^{2}).\end{split}$ (81)
@@
$\begin{split}\operatorname{df}(\hat{f}_{k,p})&=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\right]\\ &=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A})\right]+\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]\\ &=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\right]-\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]+\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right].\end{split}$ (82)
@@
$\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\right]=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\boldsymbol{\varepsilon}\right]=k.$ (83)
@@
$\begin{split}\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}&=\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\boldsymbol{\varepsilon}+\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{f}\\ &\leq\|\boldsymbol{\varepsilon}\|^{2}_{2}+\sqrt{kn}\|\boldsymbol{\varepsilon}\|_{2}.\end{split}$ (84)
In particular, we get
$\mathbb{E}\left[(\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y})^{2}\right]^{1/2}\leq C(\sqrt{k}+\sigma)\sigma n,$ (85)
@@
$\begin{split}\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]&\leq\frac{1}{\sigma^{2}}\mathbb{E}\left[(\boldsymbol{\varepsilon}^{\top}\mathbf{H}\mathbf{y})^{2}\right]^{1/2}\mathbb{P}\left(\mathcal{A}\right)^{1/2}\\ &\leq C(\sqrt{k}/\sigma+1)nk(p-k)(\pi\sigma^{2}/n)^{-1/4}\exp(-n/8\sigma^{2}).\end{split}$ (86)
@@
$\operatorname{df}(\hat{f}_{k+1,p})-\operatorname{df}(\hat{f}_{k,p})=\operatorname{df}(\hat{\beta}_{\tau}\psi_{\tau}).$ (88)
@@
$\begin{split}\operatorname{df}(\hat{\beta}_{\tau}\psi_{\tau})&=\frac{1}{\sigma^{2}}\mathbb{E}\left[(\mathbf{y}-\mathbb{E}y)^{\top}\hat{\beta}_{\tau,p}\mathbf{x}_{\cdot j,p}\right]\\ &=\frac{n}{\sigma^{2}}\mathbb{E}\left[\hat{\beta}_{\tau,p}^{2}\right]\\ &\geq\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\mathbb{1}(\mathcal{A})\right]\\ &=\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\right]-\mathbb{E}\left[\sum_{j=k+1}^{p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\mathbb{1}(\mathcal{A}^{c})\right].\end{split}$ (90)
Since $n\hat{\beta}_{j,p}^{2}/\sigma^{2}\sim_{\text{IID}}\chi^{2}_{1}$ for $k+1\leq j\leq p$, Lemma A.6 implies that the first term is lower bounded by $\frac{1}{2}\log(p-k)$ for $p-k$ large enough. Applying Cauchy-Schwarz to the second term, we can bound it (in absolute value) from above by
$(p(p+2))^{1/2}(\pi\sigma^{2}/n)^{-1/4}\exp(-n/8\sigma^{2}).$ (91)
@@
$\operatorname{df}(\hat{f}_{k+1,p})-\operatorname{df}(\hat{f}_{k,p})\geq\frac{1}{2}\log(p-k)+O(e^{-cn}).$ (92)
@@
$\operatorname{df}(\hat{f}_{k,m,p})-\operatorname{df}(\hat{f}_{k,p})\geq\frac{1}{2}(w_{k}^{k,m,p}-w_{k+1}^{k,m,p})\log(p-k)-k+O(e^{-cn}).$ (94)
Taking a limit as $n,m,p$ go to infinity finishes the proof. ∎
Lemma A.6.
@@
$\mathbb{E}\left[\max_{1\leq i\leq N}Z_{i}\right]\geq\frac{1}{2}\log N.$ (95)

### v3

@@ Acknowledgements
$\alpha_{p}^{\star}=\log\Big(\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\cdot\frac{A}{B}\Big),$ (S.43)
@@
$\alpha_{p}^{\star}=\log\Big(\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\cdot\frac{A}{B}\Big)\geq\log\Big(\frac{4}{3}\cdot\frac{1+e^{-\alpha_{0}^{\star}}}{1-e^{-\alpha_{0}^{\star}}}\Big)=\alpha_{0}^{\star}.$ (S.44)
Thus, substituting $\alpha_{p}^{\star}$ and $\alpha_{0}^{\star}$ into (S.42) and taking $\epsilon=0.05$ gives us the lower bound on expected training error gap (S.21) as
@@
$\mathbb{E}\Big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\min_{m}\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}\Big]$ $\geq\frac{1}{4}\frac{\big(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big)^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}.$
S.1.4 Proofs for Section 4.2
@@
We claim that whenever $m^{\prime}>m$, the sequence $\big\{w_{j}^{k,m^{\prime},p}\big\}_{j=1}^{p}$ majorizes the sequence $\big\{w_{j}^{k,m,p}\big\}_{j=1}^{p}$. This means that for each $l\geq 1$,
$\sum_{j=1}^{l}w_{j}^{k,m^{\prime},p}\geq\sum_{j=1}^{l}w_{j}^{k,m,p},$ (S.45)
@@
Assuming the majorization in (S.45) holds, the theorem follows directly from Theorem A.3 in Chapter 5 from marshall11 using, in their notation, $g_{i}(z)=(\operatorname{df}(\hat{f}_{i,p})-\operatorname{df}(\hat{f}_{i-1,p}))z$.
We return to proving the claim. Let $S_{k,m,p}$ denote the indices of the features selected by one inner loop of EFS with parameters $k$, $m$, and $p$. We may write
$\begin{split}\sum_{j=1}^{l}w_{j}^{k,m,p}&=\sum_{j=1}^{l}\mathbb{P}(j\in S_{k,m,p})\\ &=\mathbb{E}\left[\sum_{j=1}^{l}\mathbb{1}(j\in S_{k,m,p})\right]\\ &=\mathbb{E}\left[Z_{k,m,p}(l)\right],\end{split}$ (S.46)
where $Z_{k,m,p}(l)=|S_{k,m,p}\cap[l]|$. Evidently, (S.45) will be established if we can produce a coupling between $S_{k,m,p}$ and $S_{k,m^{\prime},p}$ such that $Z_{k,m^{\prime}p}(l)\geq Z_{k,m,p}(l)$ for $l=1,2,\ldots,p$.
[…]
We now show the desired property inductively in $k$, with the base case $k=0$ trivial. Assume the inductive hypothesis. Let $b=\min B_{m}$. Then one can show that the index selected at iteration $k+1$ (with parameter $m$) is equal to $l_{*}=\min\left\{l\colon b\leq l-Z_{k,m,p}(l)\right\}$. We therefore have
$Z_{k+1,m,p}(l)=\begin{cases}Z_{k,m,p}(l)&\text{if}~l<l_{*}\\ Z_{k,m,p}(l)+1&\text{if}~l\geq l_{*}.\end{cases}$ (S.47)
Similarly, denoting $b^{\prime}=\min B_{m^{\prime}}$ and letting $l_{*}^{\prime}=\min\left\{l\colon b^{\prime}\leq l-Z_{k,m^{\prime},p}(l)\right\}$, we get
$Z_{k+1,m^{\prime},p}(l)=\begin{cases}Z_{k,m^{\prime},p}(l)&\text{if}~l<l_{*}^{\prime}\\ Z_{k,m^{\prime},p}(l)+1&\text{if}~l\geq l_{*}^{\prime}.\end{cases}$ (S.48)
To compare $Z_{k+1,m^{\prime},p}$ and $Z_{k,m,p}$, we further set $\tilde{l}=\min\left\{l\colon b\leq l-Z_{k,m^{\prime},p}(l)\right\}$ and construct the function $\tilde{Z}$ via
[…]
@@
[…]
@@
$Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)=(Z_{k,m^{\prime},p}(l)-l)-(Z_{k,m,p}(l)-l).$ (S.51)
Because these functions take integral values, the only way $Z_{k,m^{\prime},p}(l)-Z_{k,m,p}(l)-1$ can be negative is if the right hand side of (S.51) is equal to 0. However, by the way in which $l_{*}$ is defined, we would have
$l-Z_{k,m^{\prime},p}(l)=l-Z_{k,m,p}(l)\geq b,$ (S.52)
@@
Utilizing the majorization in (S.45), we now turn to prove Theorem 5.1.
@@
$\hat{\beta}_{i,p}-\hat{\beta}_{j,p}=(\beta_{i,p}-\beta_{j,p})+\langle\mathbf{x}_{\cdot i,p}-\mathbf{x}_{\cdot j,p},\boldsymbol{\varepsilon}\rangle.$ (S.53)
@@
$\mathbb{P}\left(\hat{\beta}_{i}-\hat{\beta}_{j}<0\right)\leq\exp(-n/4\sigma^{2}).$ (S.54)
@@
$\mathbb{P}\left(\hat{\beta}_{i}+\hat{\beta}_{j}<0\right)\leq\exp(-n/4\sigma^{2}).$ (S.55)
@@
$\begin{split}\mathbb{P}\left(|\hat{\beta}_{j,p}|>|\hat{\beta}_{i,p}|\right)&\leq\mathbb{P}\left(|\hat{\beta}_{j,p}|>\hat{\beta}_{i,p}\right)\\ &\leq\mathbb{P}\left(\hat{\beta}_{j,p}>\hat{\beta}_{i,p}\right)+\mathbb{P}\left(-\hat{\beta}_{j,p}<\hat{\beta}_{i,p}\right)\\ &\leq 2\exp(-n/4\sigma^{2}).\end{split}$ (S.56)
Let $\mathcal{A}$ be the event on which $\operatorname{rank}(|\hat{\beta}_{j,p}|)\leq k$, for $j=1,\ldots,k$. We have
$\begin{split}\mathbb{P}\left(\mathcal{A}\right)&=\mathbb{P}\left(|\hat{\beta}_{i,p}|>|\hat{\beta}_{j,p}|,1\leq i\leq k<j\leq p\right)\\ &\geq 1-2k(p-k)\exp(-n/4\sigma^{2}).\end{split}$ (S.57)
@@
$\begin{split}\operatorname{df}(\hat{f}_{k,p})&=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\right]\\ &=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A})\right]+\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]\\ &=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\right]-\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]+\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right].\end{split}$ (S.58)
@@
$\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\mathbf{y}\right]=\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\widetilde{\mathbf{H}}_{k}\boldsymbol{\varepsilon}\right]=k.$ (S.59)
@@
$\begin{split}\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y}&=\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\boldsymbol{\varepsilon}+\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{f}\\ &\leq\|\boldsymbol{\varepsilon}\|^{2}_{2}+\|\boldsymbol{\varepsilon}\|_{2}\|\mathbf{f}\|_{2}.\end{split}$ (S.60)
Using the fact that $\|\mathbf{f}\|_{2}^{2}=kn$, we get
$\mathbb{E}\left[(\boldsymbol{\varepsilon}^{\top}\mathbf{H}_{k}\mathbf{y})^{2}\right]^{1/2}\leq C(\sqrt{k}+\sigma)\sigma n,$ (S.61)
@@
$\begin{split}\frac{1}{\sigma^{2}}\mathbb{E}\left[\boldsymbol{\varepsilon}^{\top}\mathbf{H}\mathbf{y}\cdot\mathbb{1}(\mathcal{A}^{c})\right]&\leq\frac{1}{\sigma^{2}}\mathbb{E}\left[(\boldsymbol{\varepsilon}^{\top}\mathbf{H}\mathbf{y})^{2}\right]^{1/2}\mathbb{P}\left(\mathcal{A}\right)^{1/2}\\ &\leq C(\sqrt{k}/\sigma+1)nk(p-k)\exp(-n/8\sigma^{2}).\end{split}$ (S.62)
@@
$\operatorname{df}(\hat{f}_{k+1,p})-\operatorname{df}(\hat{f}_{k,p})=\operatorname{df}(\hat{\beta}_{\tau}\psi_{\tau}).$ (S.64)
@@
$\begin{split}\operatorname{df}(\hat{\beta}_{\tau}\psi_{\tau})&=\frac{1}{\sigma^{2}}\mathbb{E}\left[(\mathbf{y}-\mathbb{E}y)^{\top}\hat{\beta}_{\tau,p}\mathbf{x}_{\cdot j,p}\right]\\ &=\frac{n}{\sigma^{2}}\mathbb{E}\left[\hat{\beta}_{\tau,p}^{2}\right]\\ &\geq\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\mathbb{1}(\mathcal{A})\right]\\ &=\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\right]-\mathbb{E}\left[\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\mathbb{1}(\mathcal{A}^{c})\right].\end{split}$ (S.66)
Since $n\hat{\beta}_{j,p}^{2}/\sigma^{2}\sim_{\text{IID}}\chi^{2}_{1}$ for $k+1\leq j\leq p$, Lemma S.1.6 implies that the first term is lower bounded by $\frac{1}{2}\log(p-k)$ for pkp-k large enough. For the second term, notice that
$\operatorname*{arg\,max}_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}\leq\sum_{k+1\leq j\leq p}\frac{n}{\sigma^{2}}\hat{\beta}_{j,p}^{2}=\frac{1}{\sigma^{2}}\boldsymbol{\varepsilon}^{\top}\left(\sum_{j=k+1}^{p}\mathbf{x}_{\cdot j,p}\mathbf{x}_{\cdot j,p}^{\top}\right)\boldsymbol{\varepsilon}\leq\frac{\|\boldsymbol{\varepsilon}\|_{2}^{2}}{\sigma^{2}}.$ (S.67)
Applying Cauchy-Schwarz to the second term, we can bound it (in absolute value) from above by
$Cn\exp(-n/8\sigma^{2}).$ (S.68)
@@
$\operatorname{df}(\hat{f}_{k+1,p})-\operatorname{df}(\hat{f}_{k,p})\geq\frac{1}{2}\log(p-k)+Cn\exp(-n/8\sigma^{2}).$ (S.69)
@@
$\operatorname{df}(\hat{f}_{k,m,p})-\operatorname{df}(\hat{f}_{k,p})\geq\frac{1}{2}(w_{k}^{k,m,p}-w_{k+1}^{k,m,p})\log(p-k)-k+C(\sqrt{k}/\sigma+1)nk(p-k)\exp(-n/8\sigma^{2}).$ (S.71)
Noting that Assumption 3.1 implies $k\leq p\leq n$ and taking a limit as $n,m,p$ go to infinity finishes the proof. ∎
Lemma S.1.6.
@@
$\mathbb{E}\left[\max_{1\leq i\leq N}Z_{i}\right]\geq\frac{1}{2}\log N.$ (S.72)